OpenStudy (anonymous):

a 10 kg crate slides down a slope that is 3m high and 5 m long with an initial velocity of 4m/s and then slides onto a horizontal surface coming to a stop 5 m after the end of the slope. what is the minimum coefficient of kinetic friction that is required on the horizontal surface to bring the crate to a stop.

OpenStudy (anonymous):



OpenStudy (anonymous):

vf at top is 4m/s so then vf2=vi2+2aΛx so vf at the bottom is 8.65 m/s therefore the deceleration of the object is also vf2=vi2+2aΛx and the deceleration is -0.865 if F = m x a then the force acting on the crate is F = 10 x -0.865 F = -8.65 fk=ukFn so uk=??? uk = 0.8826. but this is wrong as the answers i have in the multiple choice give 1)0.3 2)0.32 3)0.6 4)0.66

OpenStudy (anonymous):

am i doing this right?

OpenStudy (anonymous):

does it say that the slide is frictionless?

OpenStudy (anonymous):

yes the only friction is on the bottom horizontal surface

OpenStudy (anonymous):

lol i'm stuck too hha

OpenStudy (anonymous):

wait 0.76 is the last answer. and then its right. how did you get there? it obviously ran out of typing space and didnt show that last answer.

OpenStudy (anonymous):

i hate the fact that these things have a character limit

OpenStudy (anonymous):

alright. hold on...

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

I'll start with v=8.65 at the end of the slope, what u got was correct. I have two methods in mind, the easier one being Work=Change in Kinetic Energy. Here is it: Work done by friction is\[(friction force ) (distance)=-\frac{ 1 }{ 2 }(mass)(8.65m/s)^2\] \[-(10kg)(9.8m/s^2)(\mu _{k})(5m)=-\frac{ 1 }{ 2 }(10kg)(8.65m/s)^2\] friction force is negative since it's opposite the motion. Solving, well get uk=0.76..

OpenStudy (anonymous):

oh wow thanks man i apreciate it. how is the method i used wrong tho thats what i dont understand.

OpenStudy (anonymous):

The second method is using\[\Delta x = \frac{ -v^2 }{ 2a }\] find a and \[a=\frac{ -v^2 }{ \Delta x }\] Since Force=(Mass)(Acceleration) \[(mg \mu _{k})=(m)(a)\] \[\mu _{k}=\frac{ -v^2 }{ g(\Delta x) }\] Solve and you'll get the same answer. :)

OpenStudy (anonymous):

thanks for that one man. I am now certain of what I was uncertain of before! thank you!

OpenStudy (anonymous):

You're welcome. I'm glad to help :)

OpenStudy (anonymous):

also you wouldnt happen to know what the force responsible for holding a car in a frictionless banked curve is?

OpenStudy (anonymous):

is it the horizontal componet of the cars weight?

OpenStudy (anonymous):

Hmm.. I'll say it's the normal force the curve exerts on the car.

OpenStudy (anonymous):

so its the reaction force to the cars weight?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

thanks for that one as well, unfortunatly i cant give u two medals otherwise i would.

OpenStudy (anonymous):

The car is moving right? It's also sort of a centripetal force.

OpenStudy (anonymous):

yea as it is going around the curve